OpenStudy (anonymous):

A rectangular pyramid has a volume of seven hundred fifty-six cubic feet. If the base has a length of 9 feet and a width of fourteen feet, what is the height of the pyramid? 1 point five feet 6 feet twelve point five feet eighteen feet

OpenStudy (anonymous):

V = 1/3lwh 756 = (1/3)(9)(14)h 756 = 42h h =

OpenStudy (anonymous):

So would it be 12.5 feet?

OpenStudy (anonymous):

@ღChloeღ

OpenStudy (anonymous):

no its would be 18ft

OpenStudy (anonymous):

V = 1/3lwh 756 = (1/3)(9)(14)h 756 = 42h h = 18 feet